OpenStudy (anonymous):

what value must replace the "?" for this polynomial to be expressed as a difference of squares x^2+"?"x-25

OpenStudy (welshfella):

difference of 2 squares is a^2 - b^2 - there's no term in x

OpenStudy (welshfella):

so ? will equal 0

OpenStudy (welshfella):

= x^2+0x-25 = x^2 - 25